# Addressing Polarization and Unfairness in Performative Prediction

### v1

@@ 7 Conclusions & Societal Impacts
Our work reveals the unfairness issues of the PS solutions of PP. We then propose novel fairness-aware algorithms to find Fair-PS solutions with the convergence holds under mild assumptions, which will facilitate trustworthy machine learning since PP has broad applications on human-related decisions ranging from college admission and loan approval to hiring practice. Meaningful future directions include relaxing the conditions for the convergence of fairness mechanisms or expanding the mechanisms to deep learning settings.
@@ Appendix A Related Works
Performative Prediction. Proposed by Perdomo et al. [4], Performative prediction focuses on machine learning problems where the model deployments influence the data distribution, and the common application scenarios include strategic classification [1, 25, 26, 27] and predictive policing [28]. Perdomo et al. [4] formulated the problem and proposed repeated risk minimization (RRM) to ensure the convergence of model parameters to a stable point under a set of sufficient and necessary conditions. Mendler-Dünner et al. [29] presented convergence results on stochastic performative optimization. Subsequent work either strived to relax the convergence condition [30, 31] or develop methods to ensure the convergence to performative optimal points under special cases [32, 33]. Brown et al. [5] extended performative prediction to stateful settings where the current data distribution is determined by both the previous states and the deployed model. For more related works on PP, we refer to this survey [34].
Long-term fairness. There is a large line of work on long-term fairness in machine learning, where we mainly refer to distributionally robust optimization (DRO) [12, 13] and soft fair regularizers [35, 36, 37]. DRO refers to the optimization problem where the loss is minimized on the worst-case distribution around current data distribution [12, 13], while soft fair regularizers are often used in traditional fair optimization frameworks to penalize unfairness. Only a few pieces of literature touched on fairness problems under performative prediction settings. [38, 39, 17, 40] focused on long-term fairness settings under strategic classification similar to Example 3.2. Zezulka and Genin [14] pointed out fairness issues in performative prediction and formulated it within causal graphs. Hu and Zhang [15] also used causality-based models to formulate and ensure short-term and long-term fairness in performative prediction, incurring strong assumptions and expensive computational costs. Raab et al. [8] studied a special case where only the agents’ retention rates are performative. Peet-Pare et al. [21] proposed to use distributionally robust optimization (DRO) to promote fairness in performative prediction, but they proposed assumptions that are hard to verify and can not be derived from other commonly accepted assumptions in PP.
@@ B.2 Venn Diagram of Concepts in PP
Previous works that study long-term fairness [2, 40] mostly assume there to be model-dependent retention rate/population fraction $p_{s}$ changes, which can all be formulated as PRD problems. Recommendation systems are typical application scenarios of SDPP and PRD since the population fraction in the next iteration depends on both the current population and the current model. The better recommendation model captures a certain user group’s interest, the more likely this user group will have high retention rate.
@@
In this paper, PDS specifically means in each group, the data distribution $\mathcal{D}_{s}$ shifts with the decision models, e.g., in strategic classification, users strategically manipulate their features for better decision outcomes. In recommendation systems, where creators strategically manipulate their features for better decision outcomes from the model [1], and users’ interests can be shifted by the recommended items.
@@
(Partially observable Markov decision process (POMDP) as PDS [40]) The population fractions of the two groups $p_{s}$ are static, i.e., $\mathcal{D}^{(t)}=\sum_{s}p_{s}\mathcal{D}_{s}^{(t)}$, but the feature distribution changes with the deployed decision model, and such changes follow a POMDP as shown in Figure 6.
@@ B.3 Fair-RERM
$t=0$, $\mathcal{D}^{(0)}$, $\rho$, $\boldsymbol{\theta}^{(0)}$, $\delta_{\theta}$, choose fair mechanism
@@
[…]
@@ B.4 General guidance of designing fairness objectives under the SDPP setting
The key to designing a fairness objective while ensuring the convergence of the iterative algorithm is to ensure the convexity of the fairness penalty. Example 4.2 shows that group loss variance cannot preserve the convexity of $\mathcal{L}_{\textsf{fair}}$, so it fails to be a plausible objective. This implies that each penalty term in terms of $P=g(\mathcal{L})$ where the convexity of $P$ is ensured can be a fairness objective. However, it can still be highly non-trivial to design a fair objective in this form due to the following reasons: (i) it is hard to design an objective as interpretable as the ones we mentioned in Section 4. The squared loss as a penalty term is well-motivated and directly penalizes the groups/samples with the higher losses; (ii) It becomes more difficult to obtain the concrete parameters for the convergence guarantee. For example, $\bar{\ell}$ in Section 5 becomes the upper bound of $\frac{\partial g}{\partial\ell}$ and this may be hard to obtain and interpret. Overall, we see the formulations in Section 4 as more appropriate choices.
@@ B.5 Only Measure Fairness at PS Solution
(Fairness violation due to performative shifts) Consider a binary strategic classification problem with two groups $a,b$ and is a threshold model $h_{\theta}(X)=\mathbbm{1}\{X\geq\theta\}$. $p_{a},p_{b}$ are fixed and both groups satisfy $\mathbb{P}(Y^{(t)}=1)=\min\{0,\max\{X^{(t)},1\}\}$, and the transitions satisfy $X^{(t)}=X^{(0)}+(0.1+0.1\cdot\boldsymbol{1}\{s=a\})\theta^{(t)}$ but do not result in any changes in labels. If at $t=0$, both groups have features $X^{(0)}$ subject to uniform distribution on $[0,1]$, then any $\theta=0.5$ satisfies accuracy parity. However, at $t=1$, group $a$ has a shifted distribution, resulting in $\theta=0.5$ to be unfair.
@@ Appendix C Additional Experiments
All experiments are run on the CPU of a Macbook Pro with Apple M1 Pro chip and 16GB memory. We use Python 3.9 for all tasks. For datasets, Yann LeCun and Corinna Cortes hold the copyright of MNIST, which is a derivative work from original NIST datasets. MNIST dataset is made available under the terms of the Creative Commons Attribution-Share Alike 3.0 license; the credit dataset has been published by Kaggle [23]; In Appendix C.1, we visualize the convergence of performative loss for all the above experiments. In Appendix C.3, we perform experiments on a performative Gaussian data classification task. In Appendix C.2, we conduct experiments on 3 groups. In Appendix C.4, we visualize the unconvergence of the fairness penalty using group loss variance as a complementary to Example 4.2; Finally, we examine the k-delayed schema on all our fairness mechanisms on Credit data in Appendix C.5 and obtain the similar results.
C.1 Model Convergence in Main Experiments
In this section, we provide visualizations of the convergence of performative loss for all experiments of the main paper in Fig. 12. Importantly, Fig. 12(d) demonstrates the unconvergence of Fair-RERM-RW and Fair-RERM-GLP when $\rho=0.3$.
C.2 Gaussian Mean Estimation with Multiple Groups
In this section, we present the experimental results on the Gaussian mean estimation task with 3 groups $s\in\{a,b,c\}$ where $p_{a}^{(0)}=0.15,p_{b}^{(0)}=0.25,p_{c}^{(0)}=0.6$ with different target values. We let $y_{a}=0.3+\epsilon,y_{b}=0.5+\epsilon,y_{c}=0.7+\epsilon$ where $\epsilon\sim\mathcal{N}(0,0.05)$. At the beginning, the Group fraction $p_{a}^{(0)}=0.3,p_{b}^{(0)}=0.7$, and the minimum group fraction for each group is $p^{min}=\{0.1,0.1\}$. At each time $t$, denote $\mathcal{R}(s,t)=\left(1-\sum_{s^{\prime}\in\{a,b\}}p^{min}_{s^{\prime}}\right)\times\frac{1}{2}\left(p_{s}^{t}+\frac{\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{i_{s}}^{(t)})}{\sum_{s^{\prime}}\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s^{\prime}}^{(t)})}\right)+p_{s}^{min}$. $-s$ is the group other than $s$, while $i_{s}$ is found by first ranking all group-wise losses in ascending order and then selecting the $(n+1-i)-th$ largest loss if group $s$ has the $i-th$ largest loss. This retention dynamic is a direct generalization of the two-group case. Next, we show the retention of group $p^{(t+1)}_{s}$ at time $t+1$ is shown as Eqn. (13), resulting in the new data distribution $D^{(t+1)}$.
C.3 Performative Gaussian data classification.
We use a synthetic dataset consisting of $1000$ samples from 2 demographic groups $s\in\{a,b\}$ with features $X=\{X_{1},X_{2}\}$ and labels $Y\in\{0,1\}$. For group $a$, $Y=\mathbbm{1}\{x_{1}-0.5x_{2}\geq 0.5\}$. For group $b$, $Y=\mathbbm{1}\{0.5x_{1}+0.5x_{2}\geq 0.5\}$. The initial group fraction and the retention mapping are the same as in the previous experiment. Using a logistic classification model for risk minimization, we first verify the convergence of performative loss in Fig. 9 and the effectiveness of Fair-RERM-RW and Fair-RERM in Fig.13(a) and Fig. 13(b). All plots demonstrate the same trends as the previous experiment.
C.4 Gaussian Mean Estimation Using Group Loss Variance
In Example 4.2, we already prove that group loss variance as a fair penalty term violates the strong convexity assumption. In this section, we show its unconvergence while performing the Gaussian mean estimation task under the same setting mentioned in Section 6 when $\rho=0.5$. The dynamics of $\|\theta_{t}-\theta_{t-1}\|$ as well as performative loss are shown as follows. Fig. 11 shows the unconvergence, which is contradicted to Fig. 12(a) where we show the convergence of Fair-RERM.
C.5 K-delayed RERM on Credit Data
We examine the effectiveness of k-delayed RERM schema on Credit Data [23] which is the same dataset used by Brown et al. [5]. Specifically, we use $k=3$ and $\epsilon=1$ to form equal-sized agent groups uniformly in the whole population to let each subgroup best respond with different speed (Example 1 in Brown et al. [5]). We also let the retention dynamic be k-delayed, adding additional challenges compared to the previous work. We examine $\rho\in\{0.3,0.7\}$ while all other settings are the same as the Credit Data experiment in Section 6.
@@ C.5 K-delayed RERM on Credit Data
Figure 11: Gaussian mean estimation using group loss variance

### v2

@@ 7 Conclusions & limitations
Our work reveals unfairness issues of the PS solutions of PP. We propose novel fairness-aware algorithms to find Fair-PS solutions with the convergence holds under mild assumptions to facilitate trustworthy machine learning. However, the theory of this work does not cover non-convex settings and it still remains an open question for the future research.
Acknowledgements
This work was funded in part by the National Science Foundation under award number IIS2202699 and IIS-2416895.
@@ Appendix A Related works
Performative Prediction. Proposed by perdomo_performative_2021, Performative prediction focuses on machine learning problems where the model deployments influence the data distribution, and the common application scenarios include strategic classification (hardt_strategic_2015; raab2021unintended; xie2024algorithmic; xie2024learning; xie2025sprint; xie2024automating; xie2024non) and predictive policing (ensign2018runaway). perdomo_performative_2021 formulated the problem and proposed repeated risk minimization (RRM) to ensure the convergence of model parameters to a stable point under a set of sufficient and necessary conditions. Mendler2020Stochastic presented convergence results on stochastic performative optimization. Subsequent work either strived to relax the convergence condition (mofakhami2023performative; zhao2022optimizing) or develop methods to ensure the convergence to performative optimal points under special cases (miller_outside_2021; izzo2021performative). brown2020performative extended performative prediction to stateful settings where the current data distribution is determined by both the previous states and the deployed model. For more related works on PP, we refer to this survey (hardt2023performative).
Long-term fairness. There is a large line of work on long-term fairness in machine learning, where we mainly refer to distributionally robust optimization (DRO) (duchi2018learning; duchi2023distributionally) and soft fair regularizers (Kami2011; zafar2017fairness; guldogan2022equal). DRO refers to the optimization problem where the loss is minimized on the worst-case distribution around current data distribution (duchi2018learning; duchi2023distributionally), while soft fair regularizers are often used in traditional fair optimization frameworks to penalize unfairness. Only a few pieces of literature touched on fairness problems under performative prediction settings. (milli2018social; humanip; Zhang_2022_ICML; Zhang_2020_Long_term) focused on long-term fairness settings under strategic classification similar to Example D.2. zezulka2023performativity pointed out fairness issues in performative prediction and formulated it within causal graphs. hu2022achieving also used causality-based models to formulate and ensure short-term and long-term fairness in performative prediction, incurring strong assumptions and expensive computational costs. Raab_Boczar_Fazel_Liu_2024 studied a special case where only the agents’ retention rates are performative. peetpare2023long proposed to use distributionally robust optimization (DRO) to promote fairness in performative prediction, but they proposed assumptions that are hard to verify and can not be derived from other commonly accepted assumptions in PP. Another contemporary work (jia2024distributionally) developed more practical assumptions for DRO under PP settings, but they do not consider SDPP settings and did not discuss the fairness issues under PP settings. xue2024distributionally applied DRO to find the PO solutions, but the algorithm convergence needs to be established in a case-by-case basis and with careful calibration. Meanwhile, they also did not discuss the fairness issues.
Finally, it may be worthwhile noting that many works of machine learning fairness exist, but most of them only consider the population as a static distribution (e.g., martinez2020minimax). We refer to caton2024fairness as a most recent comprehensive survey.
@@ B.1 Notation Table
$d_{\boldsymbol{\theta}}$ The fixed point distribution by repeatedly deploying $\boldsymbol{\theta}$
@@ B.2 Venn diagram of concepts in PP
Previous works that study long-term fairness (Zhang_2019_Retention; Zhang_2020_Long_term) mostly assume there to be model-dependent retention rate/population fraction $p_{s}$ changes, which can all be formulated as PRD problems. Recommendation systems are typical application scenarios of SDPP and PRD since the population fraction in the next iteration depends on both the current population and the current model. The better recommendation model captures a certain user group’s interest, the more likely this user group will have high retention rate.
@@
In this paper, PDS specifically means in each group, the data distribution $\mathcal{D}_{s}$ shifts with the decision models, e.g., in strategic classification, users strategically manipulate their features for better decision outcomes. In recommendation systems, where creators strategically manipulate their features for better decision outcomes from the model (hardt_strategic_2015), and users’ interests can be shifted by the recommended items.
@@
(Partially observable Markov decision process (POMDP) as PDS (Zhang_2020_Long_term)) The population fractions of the two groups $p_{s}$ are static, i.e., $\mathcal{D}^{(t)}=\sum_{s}p_{s}\mathcal{D}_{s}^{(t)}$, but the feature distribution changes with the deployed decision model, and such changes follow a POMDP as shown in Figure 5.
@@ B.3 Fair-RERM
0:$t=0$, $\mathcal{D}^{(0)}$, $\rho$, $\boldsymbol{\theta}^{(0)}$, $\delta_{\theta}$, choose fair mechanism
@@
Suppose $\exists\alpha>1,\mu>0$ such that $\int_{\mathbb{R}^{m}}e^{\mu|x|^{\alpha}}Zdx$ is finite $\forall Z\in\triangle(\mathcal{Z})$. For a given convergence radius $\delta\in(0,1)$, take $n_{t}=\mathcal{O}\left(\frac{\log(t/p)}{(\epsilon(1+\tilde{\beta}/\gamma)\delta)^{m}}\right)$ samples at $t$. If $2\epsilon(1+\tilde{\beta}/\gamma)<1$, then with probability 1p1-p, the iterates of fair-RERM are within a radius $\delta$ of the Fair-PS solution for $t\geq(1-2\epsilon(1+\tilde{\beta}/\gamma)O(\log(1/\delta))$.
@@ B.4 General guidance of designing fairness objectives under the SDPP setting
The key to designing a fairness objective while ensuring the convergence of the iterative algorithm is to ensure the convexity of the fairness penalty. Example 4.2 shows that group loss variance cannot preserve the convexity of $\mathcal{L}_{\textsf{fair}}$, so it fails to be a plausible objective. Generally, denote the penalty term as $\mathcal{P}(\boldsymbol{\theta},\mathcal{D},\rho)$ and the fair-regularized loss $\mathcal{L}_{\textsf{fair}}=\mathcal{L}+\mathcal{P}(\boldsymbol{\theta},\mathcal{D},\rho)$, we can derive a general sufficient condition for the convergence to a unique Fair-PS solution for Fair-RRM and Fair-RERM in Thm. B.4.
Theorem B.4 (Convergence of general fairness objectives).
If $\mathcal{P}(\boldsymbol{\theta},\mathcal{D},\rho)$ is convex and $\beta_{P}$-smooth, and $\tilde{\beta}=\beta_{P}+\beta$. Then same convergence results in Thm. 5.2 and Thm. B.3 apply to $\mathcal{L}_{\textsf{fair}}$ with only the different $\tilde{\beta}$ values dependent on $\mathcal{P}$.
Proof of Thm. B.4 is almost identical to Thm. 5.2 and Thm. B.3, while the only difference lies in identifying that $\mathcal{L}_{\textsf{fair}}$ is $\tilde{\beta}-$smooth.
However, it can still be highly non-trivial to design a fair objective in practice due to the following reasons: (i) it is hard to design an objective as interpretable as the ones we mentioned in Section 4. The squared loss as a penalty term is well-motivated and directly penalizes the groups/samples with the higher losses; (ii) It becomes more difficult to obtain the concrete parameters for the convergence guarantee. For example, $\bar{\ell}$ in Section 5 becomes the upper bound of $\frac{\partial g}{\partial\ell}$ and this may be hard to obtain and interpret. Overall, we see the formulations in Section 4 as more appropriate choices because loss disparity is applicable to both classification and regression settings. Also, many prominent works on recommendation systems (e.g., (hashimoto2018fairness; Zhang_2019_Retention)) assumed the well-being of users are determined by the loss.
@@ B.5 Only measure fairness at PS solution
(Fairness violation due to performative shifts) Consider a binary strategic classification problem with two groups $a,b$ and is a threshold model $h_{\theta}(X)=\mathbf{1}\{X\geq\theta\}$. $p_{a},p_{b}$ are fixed and both groups satisfy $\mathbb{P}(Y^{(t)}=1)=\min\{0,\max\{X^{(t)},1\}\}$, and the transitions satisfy $X^{(t)}=X^{(0)}+(0.1+0.1\cdot\boldsymbol{1}\{s=a\})\theta^{(t)}$ but do not result in any changes in labels. If at $t=0$, both groups have features $X^{(0)}$ subject to uniform distribution on $[0,1]$, then any $\theta=0.5$ satisfies accuracy parity. However, at $t=1$, group $a$ has a shifted distribution, resulting in $\theta=0.5$ to be unfair.
B.6 Additional corollaries on PS-fairness guarantee
Below we provide an example corollary derived from Theorem 5.4, and we can similarly derive the PS-Fairness guarantee under other choices of algorithms and mechanisms.
Corollary B.6.
In Algorithm 1, under we can obtain $\rho_{max}$ under different fairness mechanisms, such that $\epsilon(1+\tilde{\beta}_{max}/\gamma)=1$, where
[…]
Then we define
$\triangle^{\text{PS}}_{\textsf{max-fair},\mathcal{L}}:=\lim_{\delta\rightarrow 0^{+}}\triangle^{\text{PS}}_{\textsf{fair},\mathcal{L}}(\rho_{max}-\delta),$ (13)
and we can guarantee that for fairness constraint such that $\triangle^{\text{PS}}_{\textsf{fair-tolerance},\mathcal{L}}\geq\triangle^{\text{PS}}_{\textsf{max-fair},\mathcal{L}}$, we can find a $\rho<\rho_{max}$ that produce a Fair-PS solution with fairness penalty below the constraint, i.e., $\triangle^{\text{PS}}_{\textsf{fair},\mathcal{L}}(\rho)<\triangle^{\text{PS}}_{\textsf{fair-tolerance},\mathcal{L}}$
@@ Appendix C Additional experiments
All experiments are run on the CPU of a Macbook Pro with Apple M1 Pro chip and 16GB memory. We use Python 3.9 for all tasks. For datasets, Yann LeCun and Corinna Cortes hold the copyright of MNIST, which is a derivative work from original NIST datasets. MNIST dataset is made available under the terms of the Creative Commons Attribution-Share Alike 3.0 license; the credit dataset has been published by (creditdata); In App. C.1, we visualize the convergence of performative loss for all the above experiments. We also visualize the evolution of minority group fractions for them. In App. C.4, we perform experiments on a performative Gaussian data classification task. In App. C.2, we conduct experiments on 3 groups. In App. C.5, we visualize the unconvergence of the fairness penalty using group loss variance as a complementary to Example 4.2; Finally, we examine the k-delayed schema on all our fairness mechanisms on Credit data in App. C.6 and obtain the similar results.
C.1 Complementary results of main experiments
(a) Dynamics of group-wise loss disparity for MNIST data when $\rho\in\{0,0.1,0.3\}$ When $\rho=0$, the policy is just RERM.
(b) Tradeoff not necessarily exists at the PS solution. “Stars” are produced by varying $\rho$.
Figure 6: MNIST results: Fair-RERM-RW (left), Fair-RERM-GLP (middle), Fair-RERM-SLP (right).
Performative MNIST classification.
MNIST data (deng2012mnist) consists of handwritten digits images. We randomly select 1000 images as the training set with 2 groups, where the first contains digits from 0 to $4$ and the second includes digits from $5$ to $9$. We consider the same group fraction dynamics as before with $p_{s}^{\min}=0.1$ for both groups. Using a 2-layer MLP for risk minimization, we perform RERM and Fair-RERM for $50$ rounds. Fig. 6(a) verifies that Fair-RERM can still improve fairness at the PS solution. However, this set of experiments entails non-convexity because of the use of a neural network and there is no guarantee a PS solution will exist. In Fig. 6(a), slight instability exists for the black and red lines ($\rho=0.3$) of Fair-RERM-RW and Fair-RERM-GLP. Notably, Fig. 6(b) demonstrates that the tradeoff between fairness and performative loss does not necessarily exist. The Fair-PS solution may converge to a local stationary point better for both fairness and performative loss. Although it is challenging to theoretically explain this phenomenon, the results suggest that our fairness mechanisms may still be useful in practice with deep learning models.
Firstly, provide visualizations of the convergence of performative loss for all experiments of the main paper in Fig. 9. Importantly, Fig. 9(d) demonstrates the unconvergence of Fair-RERM-RW and Fair-RERM-GLP when $\rho=0.3$. Secondly, we provide the evolution of minority group fractions for all three datasets in Fig. 10. These results clearly demonstrate our fairness mechanisms are effective in relieving polarization effects. Finally, we visualize the group-wise loss at Fair-PS solutions when $\rho$ increases in Fig. 11. It is worth noting that the fairness mechanisms do not always result in the majority group having higher loss and the minority group having lower loss. It is possible that the Fair-PS solutions are better for both groups (the right plot of Credit data and the middle/right plots of MNIST data).
@@ C.1 Complementary results of main experiments
Figure 7: Dynamics of $\theta$ and performative loss in Gaussian mean estimation setting while using group loss variance as the fairness mechanism and $\rho\in\{0.5,1.0,2.0\}$. The dynamics demonstrate non-convergence of this mechanism.
(a) Credit (EO)
(b) Gaussian (EO)
(c) Credit (DP)
(d) Gaussian (DP)
Figure 8: Equal opportunity (EO) and demographic parity (DP) disparities on Credit and Gaussian classification datasets with $\rho=0.3$.
@@
Figure 10: The evolutions of minority group fractions: top (Gaussian mean estimation); middle (Credit data); bottom (MNIST data).
Figure 11: Group-wise loss at Fair-PS solutions while varying $\rho$: top (Gaussian mean estimation); middle (Credit data); bottom (MNIST data).
C.2 Gaussian mean estimation with multiple groups
In this section, we present the experimental results on the Gaussian mean estimation task with 3 groups $s\in\{a,b,c\}$ where $p_{a}^{(0)}=0.15,p_{b}^{(0)}=0.25,p_{c}^{(0)}=0.6$ with different target values. We let $y_{a}=0.3+\epsilon,y_{b}=0.5+\epsilon,y_{c}=0.7+\epsilon$ where $\epsilon\sim\mathcal{N}(0,0.05)$. The minimum group fraction for each group is $p^{min}=\{0.1,0.1,0.1\}$. At each time $t$, denote $\mathcal{R}(s,t)=\left(1-\sum_{s^{\prime}\in\{a,b\}}p^{min}_{s^{\prime}}\right)\times\frac{1}{2}\left(p_{s}^{t}+\frac{\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{i_{s}}^{(t)})}{\sum_{s^{\prime}}\mathcal{L}(\boldsymbol{\theta}^{(t)};\mathcal{D}_{s^{\prime}}^{(t)})}\right)+p_{s}^{min}$. s-s is the group other than $s$, while $i_{s}$ is found by first ranking all group-wise losses in ascending order and then selecting the (n+1i)th(n+1-i)-th largest loss if group $s$ has the ithi-th largest loss. This retention dynamic is a direct generalization of the two-group case. Next, we show the retention of group $p^{(t+1)}_{s}$ at time $t+1$ is shown as Eqn. (14), resulting in the new data distribution $D^{(t+1)}$.
C.3 ACSIncome data classification.
we conduct an additional set of experiments on the ACSIncome-CA dataset (ding2022retiringadultnewdatasets), a pre-processed version of the widely used Adult dataset from recent fairness literature. Specifically, the prediction target is whether an individual’s annual income exceeds $50,000$ USD. We use all numerical features (AGEP, SCHL, WKHP) for RERM and Fair-RERM, with AGEP serving as the sensitive attribute (Age ¿ 35). All other settings align exactly with the Credit data experiments in Section 6. We show the dynamics of group-wise loss disparity and evolution of the majority group fraction when $\rho=\{0,1.0,3.0\}$.
Figure 14: Dynamics of group-wise loss disparity for ACSIncome-CA data when $\rho\in\{0,1.0,3.0\}$. When $\rho=0$, the policy is just RERM It is clear that when $\rho$ becomes larger, the group-wise loss disparity becomes lower.
Figure 15: The evolution of majority group for ACSIncome-CA data when $\rho\in\{0,1.0,3.0\}$. When $\rho=0$, the policy is just RERM. It is clear that when $\rho$ becomes larger, the group fraction becomes lower.
C.4 Performative Gaussian data classification.
We use a synthetic dataset consisting of $1000$ samples from 2 demographic groups $s\in\{a,b\}$ with features $X=\{X_{1},X_{2}\}$ and labels $Y\in\{0,1\}$. For group $a$, $Y=\mathbf{1}\{x_{1}-0.5x_{2}\geq 0.5\}$. For group $b$, $Y=\mathbf{1}\{0.5x_{1}+0.5x_{2}\geq 0.5\}$. The initial group fraction and the retention mapping are the same as in the previous experiment. Using a logistic classification model for risk minimization, we first verify the convergence of performative loss in Fig. 17 and the effectiveness of Fair-RERM-RW and Fair-RERM in Fig.16(a) and Fig. 16(b). All plots demonstrate the same trends as the previous experiment.